### v1

# Optimal Real-Time Fusion of Time-Series Data Under Rényi Differential Privacy

Abstract
In this paper, we investigate the optimal real-time fusion of data collected by multiple sensors. In our set-up, the sensor measurements are considered to be private and are jointly correlated with an underlying process. A fusion center combines the private sensor measurements and releases its output to an honest-but-curious party, which is responsible for estimating the state of the underlying process based on the fusion center’s output. The privacy leakage incurred by the fusion policy is quantified using Rényi differential privacy. We formulate the privacy-aware fusion design as a constrained finite-horizon optimization problem, in which the fusion policy and the state estimation are jointly optimized to minimize the state estimation error subject to a total privacy budget constraint. We derive the constrained optimality conditions for the proposed optimization problem and use them to characterize the structural properties of the optimal fusion policy. Unlike classical differential privacy mechanisms, the optimal fusion policy is shown to adaptively allocates the privacy budget and regulates the adversary’s belief in a closed-loop manner. To reduce the computational burden of solving the resulting constrained optimality equations, we parameterize the fusion policy using a structured Gaussian distribution and show that the parameterized fusion policy satisfies the privacy constraint. We further develop a numerical algorithm to jointly optimize the fusion policy and state estimator. Finally, we demonstrate the effectiveness of the proposed fusion framework through a traffic density estimation case study.
I Introduction
I-A Motivation
Information fusion from multiple sensors is a fundamental technique for environmental perception, monitoring, and control. By integrating heterogeneous measurements, a fused estimate can achieve higher accuracy, robustness, and reliability than any individual sensor alone. However, multi-source data collection and fusion often involve sensitive information, which raises significant privacy concerns. For example, traffic density estimation commonly relies on vehicles’ position and velocity data that can be automatically collected by smart devices, e.g., e.g., GPS-enabled smartphones. Although such measurements are critical for real-time traffic management and congestion prediction, they may inadvertently disclose sensitive personal attributes, such as a driver’s home and workplace locations, thereby exposing private behavioral patterns. To mitigate these risks, data collected from multiple sources are typically aggregated and processed through privacy-preserving mechanisms before being utilized for downstream estimation and control tasks.
Due to its strong mathematical guarantees and ease of implementation, differential privacy (DP) has been widely adopted for privacy-preserving data disclosure. Under DP, a privacy mechanism releases perturbed data at the cost of privacy leakage and halts data sharing once a prescribed privacy budget is exhausted. In online sensor fusion, where measurements arrive sequentially, the fusion policy should adaptively regulate its behavior and appropriately allocate the privacy budget over time to ensure sustained privacy protection while maximizing estimation performance. Otherwise, the utility of the fusion result may be significantly degraded if the privacy budget is allocated improperly or if the fusion policy is poorly designed. Despite its practical importance, optimal sensor fusion with adaptive privacy budget allocation for general nonlinear systems has not yet been thoroughly investigated in the existing literature.
I-B Related Work
Privacy-preserving methods for dynamical systems have been primarily investigated from two perspectives: information-theoretic privacy and differential privacy. Information-theoretic approaches typically quantify privacy leakage using metrics such as conditional entropy and mutual information, and formulate privacy-aware estimation and control as constrained optimization problems that balance utility and privacy [1]. In [2], the authors designed a privacy filter to protect private states in hidden Markov models by optimizing the trade-off between monitoring accuracy and privacy leakage measured by the conditional entropy of the private states conditioned on the filter output. The works in [3, 1] investigated state privacy protection in optimal linear control using directed information as the privacy metric, and showed that the optimal privacy mechanism consists of a Kalman filter followed by an additive noise module.
The authors of [4] studied the optimal privacy-aware data-sharing policy for Markov systems and proposed an actor–critic algorithm for policy optimization. In [5], privacy leakage in smart metering systems was mitigated by embedding a rechargeable battery, and the optimal battery charging policy was derived using dynamic programming. The state trajectory obfuscation problem in partially observable Markov decision processes (POMDPs) was investigated in [6], where it was shown that the optimal control law remains equivalent to the standard POMDP solution, and a piecewise-linear concave approximation approach was proposed to obtain bounded-error solutions. The optimal stochastic sampling approach for protecting private inputs in dynamic systems was studied in [7], which demonstrated that the privacy-aware sampler regulates the adversary’s belief in a closed-loop manner. However, as noted in [7, 8], the evaluation of information-theoretic privacy metrics is often computationally demanding, which renders exact optimization intractable in many practical settings.
Differential privacy (DP) [9] has also been introduced into dynamical systems to mitigate privacy leakage arising from the release of sensor measurements [10]. In [10], the authors proposed a differentially private Kalman filter by injecting additive Gaussian noise to protect private inputs, which was later extended to differentially private sensor fusion for linear systems in [11]. Differentially private linear–quadratic control for multi-agent systems was investigated in [12], where Gaussian noise was added to measurements to protect agents’ states. In [13], a Laplacian consensus algorithm was developed to preserve the privacy of agents’ initial states. The work in [14] established connections between differential privacy and input observability for linear systems and proposed a privacy-preserving controller design to protect tracking errors. In [15], the authors developed a differentially private Kalman filter through measurement aggregation and Gaussian noise injection, and showed that the optimal data aggregator minimizing the estimation error can be obtained by solving a semidefinite program. Privacy-preserving traffic density estimation under a known traffic model was studied in [16], where a moving-horizon estimation framework was proposed and validated using traffic micro-simulation software. A comprehensive and rigorous survey of differential privacy for fusion and estimation in dynamical systems was recently provided in [17].
However, most existing DP mechanisms for dynamical systems rely on assumptions of linearity and Gaussian disturbances, while optimal differentially private estimation and control for general nonlinear systems remain largely open. Moreover, in many existing works, the privacy mechanism and its parameters, such as the maximum allowable per-step privacy leakage, are assumed to be time-invariant or fixed a priori when interacting with real data [13, 14, 12]. Recent studies [18, 19, 20] have shown that adaptively selecting mechanisms and budgeting can significantly improve data utility under a fixed total privacy budget in the framework of Rényi differential privacy. It was also noted in [21] that a budgeting scheme for sequential queries, adapted to the outputs of previous queries, can improve the trade-off between data utility and privacy. Despite these advances, adaptive DP mechanisms and privacy budget allocation strategies that incorporate the dynamics of general systems have received limited attention in the literature.
Figure 1: Privacy-aware sensor fusion with adaptive budget allocation
I-C Contributions
In this paper, we propose an optimal privacy-aware sensor fusion framework for general nonlinear systems, as illustrated in Fig. 1. The system state $X_{k}$ is observed by $m$ spatially distributed sensors, and the collected measurements are aggregated into the observation vector $Y_{k}=[Y_{k}(i)]_{i=1}^{m}$, where $Y_{k}(i)$ denotes the measurement from the $i$th sensor. To prevent potential privacy leakage caused by direct data disclosure, we design a privacy-aware fusion policy that releases a fused output $Z_{k}$ in an online manner, subject to the constraint that the total privacy leakage over a finite horizon $K$ does not exceed a prescribed privacy budget $\mathsf{B_{G}}$. A remote estimator then utilizes $Z_{k}$ to reconstruct the system state.
The main contribution of this work is the development of an optimal differentially private sensor fusion framework with adaptive privacy budget allocation. Specifically, the fusion policy and the state estimator are jointly designed to minimize the state estimation error under a hard constraint on the total allowable privacy leakage $\mathsf{B_{G}}$ over a finite horizon, where the privacy leakage is quantified using Rényi differential privacy (RDP). We derive the constrained optimality equations for the optimal fusion design and show that the proposed fusion policy automatically allocates the privacy budget $\mathsf{B_{G}}$ over time in a closed-loop manner. We further develop a practical numerical algorithm to implement the proposed fusion design. The fusion policy is parameterized as a structured Gaussian mechanism that first filters measurements from individual sensors and then aggregates the filtered outputs using an adaptive fusion vector. The RDP condition of the proposed fusion policy is explicitly derived and enforced numerically through value clipping. The filtering function and the state estimator are jointly optimized by minimizing the mean estimation error over sampled trajectories, while the fusion vector is optimized using a proximal policy optimization algorithm to enable adaptive privacy budget allocation.
Compared with most existing differentially private mechanisms for linear Gaussian systems, e.g., [13, 14, 12], the proposed optimal sensor fusion design not only minimizes the state estimation error under a global differential privacy constraint for general nonlinear systems, but also adaptively determines its fusion output and allocates the privacy budget based on real-time measurements and past decisions. Furthermore, the effectiveness of the proposed privacy-aware design is demonstrated using the U.S. Highway 101 traffic dataset. Experimental results on traffic flow density estimation validate that the proposed method achieves a desirable trade-off between privacy protection and estimation accuracy.
II Outline
The remainder of this paper is organized as follows. Section IV introduces the system model and formulates the optimization problem for privacy-aware fusion design. Section V investigates the structural properties of the proposed fusion design. Section VI develops a numerical algorithm to optimize the privacy-aware fusion design. Simulation results demonstrating the effectiveness of the proposed fusion policy are presented in Section VII, followed by conclusions in Section VIII.
III Notation
We use uppercase letters, e.g., $X$, $Y$, to denote random variables, and lowercase letters, e.g., $x$, $y$, to denote their realizations. The shorthand $X^{K}$ represents the sequence $X_{0},X_{1},\dots,X_{K}]$. The symbols $p(X)$ and $p(X|Y)$ denote the probability density function and the conditional probability density function, respectively. The expectation of an function $f\left(X\right)$ is denoted with $\mathsf{E}\left[f(X)\right]=\int{p\left(x\right)f(x)dx}$.
IV Problem Formulation
IV-A System Model
We consider a general nonlinear stochastic process whose state $X_{k}\in\mathbb{R}^{n_{x}}$ evolves according to the transition kernel
$X_{k}\sim p\!\left(X_{k}\middle|X_{k-1}\right).$ (1)
The process state is observed by $m$ spatially distributed sensors, whose collective measurements at time $k$ are represented by the vector
$Y_{k}=[Y_{k}(i)]_{i=1}^{m},$ (2)
where $Y_{k}(i)\in\mathbb{R}^{n_{y}}$ denotes the local observation obtained from the $i$th sensor. The measurement trajectory $\{Y_{k}\}_{k=1}^{K}$ is private and must be kept confidential. As illustrated in Fig. 1, a fusion policy aggregates the distributed sensor data and releases an instantaneous fused output $Z_{k}\in\mathbb{R}^{n_{z}}$ to a remote estimator. An adversary is assumed to infer the private measurements from the released fusion results, which leads to privacy leakage. Since fusion outputs are released sequentially, the cumulative privacy leakage increases over time. The total allowable privacy leakage over the operation horizon $K$ is referred to as the privacy budget $\mathsf{B_{G}}\in\mathbb{R}^{+}$.
Figure 2: Traffic density estimation through randomly reported position and speed measurements
IV-B Motivating Example
Traffic density on road networks is a key performance indicator for transportation monitoring and control. It can be inferred from randomly collected position and speed measurements reported by smart devices [22]. We consider a scenario in which a subset of vehicles is equipped with similar smart devices, e.g., mobile phones operating under the same operating system or map navigation software, which is common in real-world scenarios. As illustrated in Fig. 2, these devices report GPS position and speed measurements to a centralized fusion center for traffic density estimation. The proposed framework is evaluated using the open-source U.S. Highway 101 dataset [23]. As shown in Fig. 3, the traffic density $X_{k}$ of a highway segment can be accurately estimated from the position and speed measurements $Y_{k}$. While such crowd-sourced sensing enables efficient congestion monitoring, the reported data may inadvertently reveal sensitive information about individual drivers, such as movement patterns or travel habits. To mitigate privacy risks, measurements from each vehicle can be fused by a differentially private fusion sensor prior to traffic density estimation, utilized only over a finite time horizon, and discarded once the associated privacy budget is exhausted.
Figure 3: Traffic density in US-101 Highway dataset
IV-C Privacy-Aware Fusion and Estimation
At each time step $k$, given the historical measurements $Y^{k}$ and past fusion results $Z^{k-1}$, the stochastic fusion policy processes the sensor measurements according to
$Z_{k}\sim\mathcal{F}_{k}\!\left(Z_{k}\middle|Y^{k},Z^{k-1}\right),$ (3)
where $\mathcal{F}_{k}$ denotes the conditional probability distribution of the fusion output $Z_{k}$. The resulting fused output $Z_{k}$ is then utilized by a remote estimator to produce the state estimate $\tilde{X}_{k}$ according to
$\tilde{X}_{k}\sim\mathcal{E}_{k}\!\left(\tilde{X}_{k}\middle|Z^{k},\tilde{X}^{k-1}\right),$ (4)
where $\mathcal{E}_{k}$ represents the estimation policy at time $k$. For convenience, we define the fusion policy as $\mathcal{F}=\{\mathcal{F}_{k}\}_{k=1}^{K}$ and the estimation policy as $\mathcal{E}=\{\mathcal{E}_{k}\}_{k=1}^{K}$.
To prevent the adversary from inferring the private measurements from the released fusion results, the fusion policy is designed such that the total privacy leakage does not exceed the prescribed privacy budget $\mathsf{B_{G}}$.
IV-D Privacy Measure
We adopt the Rényi differential privacy (RDP) to quantify the privacy leakage of the sensitive measurements $Y^{k}$ through the fusion mechanism. At each time step $k$, given the fusion policy and past fusion results $Z^{k-1}$, the Rényi privacy loss is defined as
$L_{k}\!\left(\mathcal{F}_{k}\left(\cdot\middle|\cdot,Z^{k-1}\right),\alpha\right)=\frac{1}{\alpha-1}\log\sup_{Y^{k},\tilde{Y}^{k}}\mathsf{E}_{\mathcal{F}_{k}\!\left(Z_{k}\middle|Y^{k},Z^{k-1}\right)}\!\left[\left.\left(\frac{\mathcal{F}_{k}\!\left(Z_{k}\middle|Y^{k},Z^{k-1}\right)}{\mathcal{F}_{k}\!\left(Z_{k}\middle|\tilde{Y}^{k},Z^{k-1}\right)}\right)^{\alpha}\right|Z^{k-1}\right],$ (5)
where $Y^{k}$ and $\tilde{Y}^{k}$ denote two adjacent measurement trajectories that differ in exactly one sensor’s measurement, i.e., $Y^{k}(j)=\tilde{Y}^{k}(j)$ for all $j\neq i$ and $Y^{k}(i)\neq\tilde{Y}^{k}(i)$ for some $1\leq i\leq m$. According to [24], the instantaneous fusion mechanism $\mathcal{F}_{k}$ satisfies $\big(\alpha,L_{k}(\mathcal{F}_{k}(\cdot|\cdot,Z^{k-1}),\alpha)\big)$-RDP.
The Rényi differential privacy characterizes the maximum divergence between the output distributions of a mechanism under neighboring datasets[24]. It serves as a natural relaxation of standard differential privacy and can be converted to the $(\epsilon,\delta)$-DP framework, while enabling more accurate and tractable privacy accounting over time. Notably, to enable privacy accounting via standard composition theorems [25, 26], the privacy loss in most existing DP designs is fixed a priori, and the mechanisms cannot adaptively adjust based on the released information. In contrast, the Rényi privacy loss defined in (5) explicitly depends on the stochastic fusion outputs, which enables the fusion policy to adaptively regulate its output and allocate the privacy budget with greater flexibility.
IV-E Optimal Privacy-Aware Fusion Design
Our objective is to develop an optimal fusion strategy and the corresponding state estimator design that minimize the estimation error of system state while ensuring that the total privacy leakage does not exceed the prescribed privacy budget $\mathsf{B_{G}}$. To this end, we formulate the optimal privacy-aware fusion problem as the following constrained optimization problem,
$\min_{\mathcal{F}}\;\min_{\mathcal{E}}\;\mathsf{E}\!\left[\sum_{k=1}^{K}d\!\left(X_{k},\tilde{X}_{k}\right)\right]$ (6)
$\mathrm{s.t.}\quad$ $\sum_{k=1}^{K}L_{k}\!\left(\mathcal{F}_{k}\!\left(\cdot\middle|\cdot,Z^{k-1}\right),\alpha\right)\leq\mathsf{B_{G}},\quad\forall Z_{k}=z_{k}\in\mathbb{R}^{n_{z}},\;k=1,\ldots,K,$ (7)
where $d\!\left(X_{k},\tilde{X}_{k}\right)$ is a measurable function that quantifies the estimation error at time step $k$. The constraint (7) ensures that the cumulative privacy leakage of each instantaneous funsion mechanism $\mathcal{F}_{k}$, given by $\sum_{k=1}^{K}L_{k}\!\left(\mathcal{F}_{k}\!\left(\cdot\middle|\cdot,Z^{k-1}\right),\alpha\right)$, remains within the prescribed privacy budget $\mathsf{B_{G}}$.
According to the fully adaptive composition theorem in [19], the constraint (7) also yields trajectory-level RDP guarantees. Specifically, $\mathcal{F}=\{\mathcal{F}_{k}\}_{k=1}^{K}$ satisfies $(\alpha,\mathsf{B_{G}})$-RDP. Consequently, solving the proposed optimization problem yields an optimal fusion policy $\mathcal{F}$ that adaptively allocates the per-step budget $L_{k}$ to minimize the estimation error subject to the trajectory-level leakage budget $\mathsf{B_{G}}$.
V Structural Properties of Optimal Privacy-Aware Fusion Design
In this section, we study the structure properties of the proposed fusion design via the dynamic programming decomposition method. We derive the constrained optimality equations for the privacy-aware fusion optimization problem, and demonstrate that the optimal fusion policy adaptively allocates the privacy budget and regulates the adversary’s belief about the private measurements.
V-A Constrained Optimality Equations
Let the the collection of instantaneous fusion policies be
$\mathcal{C}_{k}\left(Z^{k-1}\right)=\left\{\mathcal{F}_{k}\left(z_{k}\middle|y^{k},Z^{k-1}\right),\forall z_{k},y_{j}\in\mathbb{R}^{n_{y}},j=1,2,\cdots k\right\},$ (8)
which includes all conditional distributions that condition on the same historical fusion results $Z^{k-1}$ but different measurements $Y^{k}$. Note that the Rényi privacy loss (5) evaluates all instantaneous fusion polices that has the same conditional information $Z^{k-1}$, i.e., the collection (8). When solving the optimization problem (6)-(7), we cannot obtain the optimal individual sensor fusion policy $\mathcal{F}_{k}\left(\cdot\middle|Y^{k}=y^{k},Z^{k-1}\right)$ when other polices $\mathcal{F}_{k}\left(\cdot\middle|Y^{k}\neq y^{k},Z^{k-1}\right)$ in the collection $\mathcal{C}_{k}$ is unknown. In other words, we need to simultaneously solve all fusion polices with same $Z^{k-1}$, i.e., optimize the policy collection $\mathcal{C}_{k}$.
We next derive the constrained Bellman equations of the policy collection $\mathcal{C}=\left\{\mathcal{C}_{k}\right\}_{k=1}^{K}$ based on the optimality principle.
Theorem 1.
The optimal differential private fusion collection $\mathcal{C}$ can be obtained via solving the following constrained optimality equations,
$J_{k}^{\star}\left(s_{k},b_{k}\right)=$ $\min_{\mathcal{C}_{k}}\mathsf{E}\left[d\left(X_{k},\tilde{X}_{k}^{\star}\right)+J_{k+1}^{\star}\left(s_{k+1},b_{k+1}\right)\middle|Z^{k-1}\right]$ (9)
[…]
where $s_{k}$ is the remaining privacy budget with $s_{k}=s_{k-1}-L_{k-1}\left(\mathcal{C}_{k-1}\left(b_{k-1},s_{k-1}\right);\alpha\right)$ and $s_{1}=\mathsf{B_{G}}$, $b_{k}\left(X_{k},Y_{k}\right)=p\left(X_{k},Y^{k}\middle|Z^{k-1}\right)$ is the belief state with the update rule $b_{k+1}=\Phi\left(b_{k},\mathcal{C}_{k},Z_{k}\right)$, i.e.,
$b_{k+1}\left(X_{k+1},Y^{k+1}\right)=\frac{p\left(Y_{k+1}\middle|X_{k+1}\right)\int{p\left(X_{k+1}\middle|x_{k}\right)p\left(x_{k},Y^{k}\middle|Z^{k-1}\right)dx_{k}}\mathcal{F}_{k}\left(Z_{k}\middle|Y^{k},Z^{k-1},\mathsf{B}_{k}\right)}{\int{\int{p\left(x_{k},y^{k}\middle|Z^{k-1}\right)\mathcal{F}_{k}\left(Z_{k}\middle|y^{k},Z^{k-1},\mathsf{B}_{k}\right)dx_{k}}dy^{k}}}.$ (13)
$J^{\star}$ is the optimal cost-to-go function with $J_{K+1}=0$. The optimal fusion collection is $\mathcal{C}^{\star}_{k}\left(b_{k},s_{k}\right)$.
Proof.
See in Appendix A. ∎
According to Theorem 1, given the released fusion results $Z^{k-1}$, we can forwardly updated the belief state $b_{k}$ and the remaining privacy $s_{k}$, and backwardly compute the optimal fusion collection by solving the constrained optimality equations (9)-(12). Also, to meet the privacy budget constraint (7), the per-step leakage $L_{k}$ through $\mathcal{C}_{k}$ cannot exceed the remaining privacy budget (12).
Base on the optimization result, given the measurements $Y^{k}$, the optimal fusion policy can be be further obtained via selecting the conditional distribution in $\mathcal{C}_{k}^{\star}$ that matches $Y^{k}$ as discussed as follows.
Lemma 1.
Given the information $\left(Y^{k},Z^{k-1}\right)$ and the optimal fusion collection $\mathcal{C}_{k}^{\star}$, the optimal fusion policy $\mathcal{F}_{k}^{\star}\left(\cdot\middle|Y^{k},s_{k},b_{k}\right)$ is obtained by selecting the conditional distribution in $\mathcal{C}_{k}^{\star}$ that corresponds to the actual measurements $Y^{k}$.
As shown in Lemma 1, the optimal fusion policy depends not only on the measurements $Y^{k}$, but also on the remaining privacy budget $s_{k}$ and the belief state $b_{k}$, which distinguishes it from classical DP fusion policies based on the standard composition theorem. The structural difference between the optimal fusion policy and classical fusion policies is illustrated in Fig. 4. As depicted in Fig. 4, the optimal fusion policy adaptively allocates the privacy budget by controlling the remaining budget $s_{k}=s_{k-1}-L_{k-1}$, while simultaneously estimating the system state via the belief state $b_{k}\left(X_{k},\cdot\mid Z^{k-1}\right)$ and controlling the adversary’s belief of the private measurements via $b_{k}\left(\cdot,Y^{k}\mid Z^{k-1}\right)$ in a closed-loop manner.
In contrast, classical fusion policies usually rely on a predetermined per-step privacy budget $\mathsf{B}=\mathsf{B_{G}}/K$ and generate fusion outputs solely based on the measurements. Consequently, such policies cannot adaptively assign more privacy budget to some important time instance to improve data utility. By incorporating both the remaining privacy budget and the belief state, the optimal fusion policy offers greater flexibility in achieving a favorable privacy–utility trade-off. This advantage is further validated through numerical experiments in Sec. VII.
Figure 4: The structure of the optimal fusion policy.
VI Numerical Algorithm
The optimal differentially private fusion design is computationally demanding, primarily due to the difficulty of evaluating the privacy leakage for general fusion policies (5) and the curse of dimensionality inherent in the dynamic programming formulation of the constrained optimality equations (9)–(12) [27]. To address these challenges, we first parameterize the fusion policy using a structured conditional Gaussian distribution and derive a simple analytical expression for the Rényi privacy loss, which simplifies privacy evaluation. We then develop a numerical algorithm to optimize the parameterized fusion policy based on proximal policy optimization.
VI-A Policy Parameterization
We adopt a parametric recurrent representation to extract historical fusion information $Z^{k-1}$,
$h_{k,z}=\mathcal{H}_{\phi,z}\left(h_{k-1,z},Z_{k-1}\right),$ (14)
where $\mathcal{H}_{\phi,z}$ is a measurable function parameterized by $\phi$. Similarly, the measurements from the $i$-th sensor, $Y^{k}(i)$, are compressed via a recurrent representation,
$h_{k,y}(i)=\mathcal{H}_{\theta,y}\left(h_{k-1,y},Y_{k}(i)\right),$ (15)
and $h_{k,y}=[h_{k,y}(i)]_{i=1}^{m}$ extracts information of measurements from different sensors.
The fusion policy is then parameterized as a conditional Gaussian distribution,
$\mathcal{F}_{\phi,\theta}\left(Z_{k}\middle|Y^{k},Z^{k-1}\right)=\mathcal{N}\Big(Z_{k};\,g_{\phi}(h_{k,z},s_{k})\,f_{\theta}(h_{k,y}),\,I\Big),$ (16)
where $f_{\theta}(h_{k,y})=[f_{\theta}(h_{k,y}(i))]_{i=1}^{m},f_{\theta}(h_{k,y}(i))\in[0,1]^{d}$ is the filtering function applied to each sensor, and $g_{\phi}(h_{k,z},s_{k})\in\mathbb{R}^{m}$ is the fusion vector. Specifically, measurements from different sensors are first processed by the filtering function $f_{\theta}$, then aggregated by the fusion vector $g_{\phi}$, and finally perturbed with Gaussian noise. The structure of the parameterized policies is illustrated in Fig. 5 for clarity.
To sample the fusion result from (16), we first draw Gaussian noise $N_{k}\sim\mathcal{N}(0,I_{d})$ from a standard Gaussian distribution, and then generate $Z_{k}$ via
$Z_{k}=g_{\phi}(h_{k,z},s_{k})f_{\theta}(h_{k,y})+N_{k}.$ (17)
For comparison, most classical differentially private mechanisms are independent of historical released information and use fixed privacy parameters, which can be similarly parameterized as
$Z_{k}=\tilde{g}f_{\theta}(h_{k,y})+N_{k},$ (18)
where the fusion vector $\tilde{g}$ is constant and determined prior to interacting with the measurements. In contrast, our fusion policy adaptively adjusts its output based not only on the current measurements but also on the historical released information, as illustrated in Fig. 5. This structure aligns with the optimal privacy-aware fusion policy described in Theorem 1. We left the further numerical comparisons in Sec. VII to highlight the benefits of using the adaptive fusion vector $g_{\phi}(h_{k,z},s_{k})$.
Furthermore, the estimation policy is parameterized as $\mathcal{E}_{\omega}(\tilde{h}_{k,z})$, with the recurrent hidden state
$\tilde{h}_{k,z}=\tilde{\mathcal{H}}_{\omega,z}(\tilde{h}_{k-1,z},Z_{k-1}).$ (19)
This parameterization enables end-to-end optimization of the fusion and estimation policies while maintaining computational tractability.
Figure 5: The structure of the parameterized fusion policy.
VI-B Privacy Leakage Evaluation
After parameterization, the privacy leakage of the proposed fusion policy (17) can be efficiently evaluated, significantly reducing the complexity of privacy computation.
Theorem 2.
Let $\bar{g}_{\phi}(h_{k,z},s_{k})=\max_{i}\left|g_{\phi}(h_{k,z},s_{k},i)\right|$. Then, the instantaneous Rényi privacy loss can be computed as
$L_{k}\left(\mathcal{F}_{k},Z^{k-1};\alpha\right)=\frac{\alpha d}{2}\,\bar{g}_{\phi}^{2}(h_{k,z},s_{k}).$ (20)
Furthermore, given the remaining privacy budget $s_{k}\geq 0$, the per-step leakage satisfies $L_{k}(\mathcal{F}_{k},Z^{k-1};\alpha)\leq s_{k}$ if
$-\sqrt{\frac{2s_{k}}{\alpha d}}\leq\bar{g}_{\phi}(h_{k,z},s_{k})\leq\sqrt{\frac{2s_{k}}{\alpha d}}.$ (21)
Proof.
See Appendix B. ∎
According to Theorem 2, the privacy leakage through $\mathcal{F}_{k}$ remains within the remaining budget as long as the fusion vector does not amplify the filtered measurements excessively relative to the Gaussian noise $N_{k}$, i.e., $\bar{g}_{\phi}$ is appropriately bounded. Notably, the constraint in (21) varies with the historical fusion results $Z^{k-1}$ encoded in $h_{k,z}$, providing the fusion policy with greater flexibility to balance estimation error (6) against privacy constraints (7). Moreover, for any fusion policy, the privacy budget constraint (12) is guaranteed if the fusion vector is clipped within the range specified by (21), i.e.,
$\text{clip}\big(g_{\phi}(h_{k,z},s_{k}),-\sqrt{\frac{2s_{k}}{\alpha d}},\sqrt{\frac{2s_{k}}{\alpha d}}\big)=\min\Big\{\max\big\{-\sqrt{\frac{2s_{k}}{\alpha d}},g_{\phi}(h_{k,z},s_{k})\big\},\sqrt{\frac{2s_{k}}{\alpha d}}\Big\}.$ (22)
VI-C $\left(\alpha,\mathsf{B_{G}}\right)$-RDP Sensor Fusion
Motivated by Theorem 2, we implement the sensor fusion algorithm (Algorithm 1) by embedding the hard constraint (12) via the clipping operation (22). Since the privacy budget constraint is never violated, Algorithm 1 satisfies $\left(\alpha,\mathsf{B_{G}}\right)$-RDP, ensuring differential privacy at the trajectory level [19].
Furthermore, Algorithm 1 can be used to evaluate the parameterized fusion policy $\mathcal{F}_{\theta,\phi}$ and the state estimator $\mathcal{E}_{\omega}$, which is necessary for applying gradient descent techniques to update their parameters. In the following, we describe in detail how to optimize the fusion policy and state estimator based on Algorithm 1.
Algorithm 1 $\left(\alpha,\mathsf{B_{G}}\right)$-RDP sensor fusion
Initialize $k=1$, $s_{1}=\mathsf{B_{G}}$, $Z_{0}=0$, $\mathsf{B}_{0}=0$, $h_{k,y}$, $h_{k,z}$ and $\tilde{h}_{k,z}$.
while $k\leq K$ do
  Collect the measurement $Y_{k}$ from the environment
  Update the recurrent representations via $h_{k,y}$, $h_{k,z}$ and $\tilde{h}_{k,z}$
  Compute the fusion vector $g_{\phi}\big(h_{k,z},s_{k}\big)$ and clip it using (22).
  Release the fusion result: $Z_{k}=g_{\phi}\left(h_{k,z},\mathsf{B}_{k}\right)f_{{\theta}}\left(h_{k,y}\right)+N_{k}$, $N_{k}\sim\mathcal{N}\left(0,I_{d}\right)$
  Compute the state estimate $\mathcal{E}_{\omega}\left(\tilde{h}_{k,z}\right)$
  Update the remaining privacy budget $s_{k+1}=s_{k}-\frac{\alpha d}{2}\bar{g}_{\phi}^{2}\left(h_{k,z},\theta\right)$
  The environment returns the estimation error $d\left(X_{k},\tilde{X}_{k}\right)$
  $k=k+1$
end while
VI-D Joint-optimization for Filtering Function and Estimator
As illustrated in Fig. 5, the measurement and fusion information propagate forward through the deterministic filtering function to the state estimator, while the fusion vector is influenced backward by past fusion results. Consequently, when the fusion vector is fixed, the filtering function $f_{\theta}$ and the state estimator $\mathcal{E}_{\omega}$ can be jointly optimized by minimizing the sample mean of the estimation error. Specifically, we first sample $M$ trajectories $\{X^{K,(i)},\tilde{X}^{K,(i)}\}_{i=1}^{M}$ according to Algorithm 1, and then update $\theta$ and $\omega$ via
$(\omega,\theta)\leftarrow\,(\omega,\theta)-\alpha\frac{1}{M}\sum_{i=1}^{M}{\sum_{k=1}^{K}{d\bigl(X_{k}^{(i)},\tilde{X}_{k}^{(i)}\bigr)}},$ (23)
where the objective (23) approximates (6) using the sample mean over the trajectories, and $\alpha$ is the step size.
VI-E Proximal Policy Optimization for Fusion Vector Function
As indicated in Theorem 1, the optimal fusion vector function $g_{\phi}$ design is a dynamic programming problem. Proximal Policy Optimization (PPO) is a robust and sample-efficient reinforcement learning approach for such dynamic programming problems. Here, we show how to optimize $g_{\phi}$ via PPO algorithm while $f_{\theta}$ and $\mathcal{E}_{\omega}$ are fixed via (23). We follow the standard PPO framework [28], defining the fusion vector as an action sampling from the stochastic policy,
$a_{k}\sim\mathcal{N}\big(a_{k};g_{\phi}(h_{k,z},s_{k}),\Sigma_{\phi}\big),$ (24)
where $\Sigma_{\phi}$ is a learnable variance controlling the exploration–exploitation trade-off. This stochastic policy is denoted $\mathcal{A}_{\phi}(a_{k}|s_{k},h_{k,z})$.
Given information $(Z^{k-1},s_{k})$, the cost-to-go function of the fusion vector policy is
$J(Z^{k-1},s_{k})=\mathsf{E}\big[Q(a_{k},Z^{k-1},s_{k})\,|\,Z^{k-1},s_{k}\big],$ (25)
where the $Q$-function evaluates the value of action $a_{k}$,
$Q(a_{k},s_{k},Z^{k-1})=\mathsf{E}\big[d(X_{k},\tilde{X}_{k})+J(s_{k+1},Z^{k})\,|\,a_{k},s_{k},Z^{k-1}\big],$ (26)
with $J(s_{K+1},Z^{K})=0$. The corresponding advantage function is
$G(a_{k},s_{k},Z^{k-1})=Q(a_{k},s_{k},Z^{k-1})-J(s_{k},Z^{k-1}),$ (27)
quantifying the relative benefit of selecting action $a_{k}$.
Following the actor–critic framework, the cost-to-go function is parameterized with $\psi$ and estimated via minimizing the mean-square error:
$\psi\leftarrow\,\psi-\beta\frac{1}{M}\sum_{i=1}^{M}{\sum_{k=1}^{K}{\bigl(J_{\psi}(s_{k}^{\left(i\right)},\hat{h}_{k,z}^{\left(i\right)})-\sum_{j=k}^{K}{d(X_{j}^{\left(i\right)},\tilde{X}_{j}^{\left(i\right)})}\bigr)^{2}}},$ (28)
with recursive representation $\hat{h}_{k,z}=\hat{\mathcal{H}}_{\psi}(\hat{h}_{k-1,z},Z_{k-1})$ and the positive step size $\beta$.
In the PPO framework, we denote the current fusion vector policy to be updated as $\mathcal{A}_{\phi}$, and the previous policy before current update as $\mathcal{A}_{\phi,\mathrm{old}}$. The probability ratio between the current and previous policies is defined as
$r_{k}(\phi)=\frac{\mathcal{A}_{\phi}(a_{k}\mid s_{k},h_{k,z})}{\mathcal{A}_{\phi,\mathrm{old}}(a_{k}\mid s_{k},h_{k,z})}.$ (29)
We then sample trajectories from Algorithm 1 with $\mathcal{A}_{\phi,\mathrm{old}}$ and compute the advantage function $G_{\mathrm{old}}(a_{k},s_{k},Z^{k-1})$, which evaluates the relative benefit of action $a_{k}$ under the previous policy.
The policy parameters $\phi$ are updated by optimizing the following surrogate objective function:
$\phi\leftarrow\phi-\gamma\frac{1}{M}\sum_{i=1}^{M}{\sum_{k=1}^{K}{\min\bigl(r_{k}(\phi)\,G_{\mathrm{old}}(a_{k}^{\left(i\right)},s_{k}^{\left(i\right)},Z^{k-1,\left(i\right)}),\;H_{\mathrm{old}}(\epsilon,G_{\mathrm{old}},a_{k}^{\left(i\right)},s_{k}^{\left(i\right)},Z^{k-1,\left(i\right)};\phi)\bigr)}},$ (30)
where the clipped term is defined as
$H_{\mathrm{old}}(\epsilon,G_{\mathrm{old}},a_{k},s_{k},Z^{k-1};\phi)=\mathrm{clip}\big(r_{k}(\phi),1-\epsilon,1+\epsilon\big)\,G_{\mathrm{old}}(a_{k},s_{k},Z^{k-1}).$ (31)
In this formulation, $r_{k}(\phi)G_{\mathrm{old}}$ measures the improvement of the current policy relative to the previous one, while the clipping operation constrains the policy ratio within $1-\epsilon,1+\epsilon]$. This ensures stable and robust policy updates, preventing overly large steps that may degrade performance.
Finally, leveraging the numerical techniques described above, Algorithm 2 is proposed to alternatively optimize the filtering function and estimator jointly, and optimize the fusion vector using PPO.
Algorithm 2 Alternative Optimization of Fusion Policy and State Estimator
Initialize the fusion function $g_{\phi}$, filtering function $f_{\theta}$ and the state estimator $\mathcal{E}_{\omega}$.
repeat
  repeat
   Collect data and evaluate the filtering function $f_{\theta}$ and state estimator $\mathcal{E}_{\omega}$ via Algorithm 1.
   Update $f_{\theta}$ and $\mathcal{E}_{\omega}$ via (23) with the fixed fusion function $g_{\phi}$.
  until convergence
  repeat
   Collect data and evaluate $g_{\phi}$ via Algorithm 1.
   Update $g_{\phi}$ via PPO algorithm with fixed $f_{\theta}$ and $\mathcal{E}_{\omega}$.
  until convergence
until convergence
VII Numerical Results
In this section, we apply the proposed privacy-aware fusion design to estimate the traffic density $X_{k}$, defined as the average number of vehicles per kilometer on a road segment, based on position and velocity data $Y_{k}$ randomly collected from vehicles.
VII-A US Highway 101 Dataset
To evaluate the robustness and effectiveness of Algorithm 2, we conduct simulations based on the US Highway 101 dataset [23], which contains 45 minutes of detailed vehicle trajectory data on southbound US 101. We select a 400-meter segment as the study area, set the sampling interval to 0.2s, and consider a horizon $K=100$.
To better reflect real-world traffic scenarios, we assume that a subset of drivers uses similar smartphones operating under the same operating systems or navigation software, and that these devices report measurements to a common fusion center for traffic prediction and management. In the simulation, a subset of vehicles entering the study area is assumed to report position and speed measurements every 0.2 s, with at most 12 vehicles transmitting simultaneously. Meanwhile, the ground-truth traffic density is computed by counting the number of vehicles within the study area at each time step and normalizing by the segment length for estimation comparison. The estimation performance of the proposed fusion scheme is evaluated using the squared error, defined as $d\left(X_{k},\tilde{X}_{k}\right)=\left(X_{k}-\tilde{X}_{k}\right)^{2}$.
The dataset is divided into training, validation, and testing subsets. Specifically, 80% of the data is used for training the fusion policy, while the remaining data is split equally, with 10% used for validation and 10% for testing.
VII-B Privacy-Aware Traffic Density Estimation
We compare the proposed adaptive privacy-aware fusion design (17) with the classical differentially private mechanism (18) that uniformly distributes the privacy budget over time. The proposed design is optimized using Algorithm 2, while the classical mechanism is optimized by minimizing the mean squared error (23).
To assess the privacy–utility trade-off, we plot the average estimation error versus the total privacy budget $\mathsf{B_{G}}$ in Fig. 6. As shown in Fig. 6, the estimation error increases as the privacy budget decreases, reflecting the trade-off between higher privacy level and lower estimation accuracy. Specifically, a smaller privacy budget limits the maximum fusion vector magnitude $\bar{g}_{\phi}$ according to (21), reducing the signal-to-noise ratio of the fused output $Z_{k}$ and increasing the estimation error.
Specifically, the proposed adaptive fusion design achieves lower estimation error than the classical mechanism under the same privacy budget by dynamically allocating the budget over time. We compare the per-step allocated budgets and the corresponding estimation errors for the two fusion designs with a total budget of $\mathsf{B_{G}}=1.5$ in Fig. 7 and Fig. 7, respectively. The adaptive design allocates more budget in the 2–8s interval, leading to lower estimation error compared to the classical design.
Moreover, we draw the estimated traffic density trajectories under the two fusion designs for $\mathsf{B_{G}}=1.5$ in Fig. 8. As observed, the trajectory obtained with adaptive budget allocation more closely matches the true traffic density, demonstrating improved estimation accuracy over the uniform budget allocation.
Figure 6: Average estimation error versus the privacy budget $\mathsf{B_{G}}$ under the adaptive fusion design and the classical fusion design.
Figure 7: (a) Privacy budget allocation, (b) estimation error under the adaptive fusion design and the classical fusion design with the budget $\mathsf{B_{G}}=1.5$.
Figure 8: Traffic density estimation with budget $\mathsf{B_{G}}=1.5$.
VIII Conclusion
In this paper, we proposed an optimal privacy-aware sensor fusion framework for general nonlinear systems using Rényi differential privacy. We analyzed the structural properties of the optimal fusion design through the derived constrained optimality equations and showed that the fusion policy adaptively allocates the privacy budget while minimizing estimation error. Furthermore, we developed a practical numerical algorithm to optimize the parameterized fusion policy and jointly train the state estimator. The effectiveness of the proposed approach was demonstrated through traffic density estimation on a real-world US Highway 101 dataset, where the adaptive privacy-aware fusion achieved improved estimation accuracy compared to classical differentially private mechanisms.
Appendix A Proof of Theorem 1
To prove Theorem 1, we first define the remaining privacy budget and belief state, and show that they can be updated in a Bayesian manner.
Lemma 2.
Let $s_{k}$ denote the remaining privacy budget with $s_{1}=\mathsf{B_{G}}$, and let $b_{k}$ denote the conditional distribution
$b_{k}=p\left(X_{k},Y^{k}\middle|Z^{k-1}\right)$
with $b_{1}(X_{1},Y_{1})=p(X_{1},Y_{1})$. Then, $s_{k+1}$ and $b_{k+1}$ can be updated as
$s_{k+1}=s_{k}-L_{k}\left(\mathcal{C}_{k}(Z^{k-1});\alpha\right),$ (32)
$b_{k+1}(X_{k+1},Y^{k+1})$ $=\frac{p(Y_{k+1}\mid X_{k+1})\int p(X_{k+1}\mid x_{k})p(x_{k},Y^{k}\mid Z^{k-1})dx_{k}\,\mathcal{F}_{k}(Z_{k}\mid Y^{k},Z^{k-1})}{\int\int p(x_{k},y^{k}\mid Z^{k-1})\mathcal{F}_{k}(Z_{k}\mid y^{k},Z^{k-1})dx_{k}dy^{k}},$ (33)
where the update of belief state $b_{k}$ is denoted as $b_{k+1}=\Phi(b_{k},Z_{k},\mathcal{C}_{k})$.
Proof.
By definition, the remaining budget is initialized as $s_{1}=\mathsf{B_{G}}$, and at each step the consumed privacy is $L_{k}(\mathcal{C}_{k}(Z^{k-1});\alpha)$, giving
$s_{k+1}=s_{k}-L_{k}(\mathcal{C}_{k}(Z^{k-1});\alpha).$
For the belief state, we have
$b_{k+1}(X_{k+1},Y^{k+1})=p(X_{k+1},Y^{k+1}\mid Z^{k})=\frac{p(X_{k+1},Y^{k+1},Z_{k}\mid Z^{k-1})}{p(Z_{k}\mid Z^{k-1})}.$
The numerator can be expanded as
$p(X_{k+1},Y^{k+1},Z_{k}\mid Z^{k-1})=\int p(Y_{k+1}\mid X_{k+1})p(X_{k+1}\mid x_{k})p(x_{k},Y^{k}\mid Z^{k-1})\mathcal{F}_{k}(Z_{k}\mid Y^{k},Z^{k-1})dx_{k},$
and the denominator is the marginal over $X_{k+1},Y^{k+1}$. This gives the update formula in the lemma. The update depends on the fusion result $Z_{k}$ and the policy collection $\mathcal{C}_{k}(Z^{k-1})$, which we summarize as $b_{k+1}=\Phi(b_{k},Z_{k},\mathcal{C}_{k})$. ∎
We now prove Theorem 1. Given a fusion policy, the optimal estimator is
$\tilde{X}_{k}^{\star}=\arg\min_{\tilde{X}_{k}}\mathsf{E}\big[d(X_{k},\tilde{X}_{k})\mid Z^{k}\big],$
since the estimator does not affect future fusion decisions or system dynamics.
The privacy constraint
$\sum_{k=1}^{K}L_{k}(\mathcal{C}(Z^{k-1});\alpha)\leq\mathsf{B_{G}}$
defines the admissible policy collection
$\mathcal{C}_{k}(Z^{k-1})\in\Big\{\mathcal{C}_{k}(Z^{k-1}):L_{k}(\mathcal{C}(Z^{k-1});\alpha)\leq s_{k}\Big\},\quad\forall k=1,\dots,K,$
with $s_{k+1}=s_{k}-L_{k}\left(\mathcal{C}_{k}(Z^{k-1});\alpha\right)$.
Base on the admissible policy collection, we define the optimal cost-to-go function at step $k$ as
$J^{\star}(Z^{k-1})=\min_{\text{admissible }\{\mathcal{C}_{j}\}_{j=k}^{K}}\mathsf{E}\Big[\sum_{j=k}^{K}d(X_{j},\tilde{X}_{j}^{\star})\mid Z^{k-1}\Big].$
By induction, assume for $k+1\leq K$ we have
$J^{\star}(Z^{k})=J^{\star}(s_{k+1},b_{k+1}),$
with $\sum_{j=k+1}^{K}L_{j}(\mathcal{C}(Z^{j-1});\alpha)\leq s_{k+1}$. Using the principle of optimality, we obtain
$J^{\star}(Z^{k-1})$ $=\min_{\text{admissible }\mathcal{C}_{k}}\mathsf{E}\big[d(X_{k},\tilde{X}_{k}^{\star})+J^{\star}(Z^{k})\mid Z^{k-1}\big]$
$=\min_{\text{admissible }\mathcal{C}_{k}}\mathsf{E}\big[d(X_{k},\tilde{X}_{k}^{\star})+J^{\star}(s_{k+1},b_{k+1})\mid Z^{k-1}\big]$
$=\min_{\text{admissible }\mathcal{C}_{k}}\mathsf{E}\big[d(X_{k},\tilde{X}_{k}^{\star})+J^{\star}(s_{k}-L(\mathcal{C}_{k}(Z^{k-1});\alpha),\Phi(b_{k},Z_{k},\mathcal{C}_{k}))\mid Z^{k-1}\big],$ (34)
where the last equality follows from Lemma 2. The expectation is taken over $p(X_{k},Z_{k}\mid Z^{k-1})$, which depends on $b_{k}$. Hence, the optimal cost-to-go depends on $s_{k}$ and $b_{k}$, and the policy collection $\mathcal{C}_{k}$ can be expressed as a function of $(s_{k},b_{k})$, yielding the constrained optimality equations in Theorem 1.
Appendix B Proof of Theorem 2
Consider adjacent measurement trajectories $Y^{k}$ and $\tilde{Y}^{k}$ that differ only at sensor $i$, i.e., $Y^{k}(i)\neq\tilde{Y}^{k}(i)$. Since each element of the filtering function processes only one sensor’s measurements, we have
$f_{\theta}(h_{k,y}(i))\neq f_{\theta}(h_{k,\tilde{y}}(i))\quad\text{and}\quad f_{\theta}(h_{k,y}(j))=f_{\theta}(h_{k,\tilde{y}}(j)),\;j\neq i.$
Thus, the $l_{2}$ sensitivity of the fusion policy (17) is
$\Delta_{2}(\mathcal{F}_{k},Z^{k-1})$ $:=\sup_{Y^{k},\tilde{Y}^{k}}\big\|g_{\phi}(h_{k,z},s_{k})\big[f_{\theta}(h_{k,y})-f_{\theta}(h_{k,\tilde{y}})\big]\big\|_{2}$
$=\max_{i}\|g_{\phi}(h_{k,z},s_{k},i)\mathsf{1}_{d}\|_{2}$
$=\max_{i}\sqrt{d}\,|g_{\phi}(h_{k,z},s_{k},i)|$
$:=\sqrt{d}\,\bar{g}_{\phi}(h_{k,z},s_{k}),$ (35)
where $\mathsf{1}_{d}$ is a $d$-dimensional vector of ones, and the second equality follows from $f_{\theta}(h_{k,y}(i))\in[0,1]^{d}$.
By [24], the Rényi privacy loss of the Gaussian mechanism is
$L_{k}(\mathcal{F}_{k}(\cdot\mid Z^{k-1}),\alpha)=\frac{\alpha}{2}\Delta_{2}^{2}(\mathcal{F}_{k},Z^{k-1})=\frac{\alpha d}{2}\bar{g}_{\phi}^{2}(h_{k,z},s_{k}).$
Consequently, if
$\sqrt{\frac{2s_{k}}{\alpha d}}\leq\bar{g}_{\phi}(h_{k,z},s_{k})\leq\sqrt{\frac{2s_{k}}{\alpha d}},$
then $L_{k}(\mathcal{F}_{k}(\cdot\mid Z^{k-1}),\alpha)\leq s_{k}$, satisfying the per-step RDP constraint.